TABLE 1           Sales         $47,000 Current assets of $ 5,100, Current liabilities      $ 6,200,    &n

TABLE 1

          Sales         $47,000

Current assets of $ 5,100,

Current liabilities      $ 6,200,

          Cost           44,650

Net fixed assets of $51,500

Owners Equity           50, 400

            

          Net Income 2,350

56,600

Owners Equ & Liab.      56,600

Sales are expected to increase by 3 percent next year. Net Income, that is, Net Profit Margin (NPM) is 5% of Sales. The firm has no long term debt and does not plan on acquiring any. The firm does not pay any dividends and all assets, short term liabilities, and costs vary directly with sales.

1. Base upon the above data in Table 1 the projected Total Asset will be]

[Work Space]

Circle A,B,C, or D

     A. $51,500

     B. $58,298

     C. $56,600

     D. Other $____

2. Based upon the above data how much additional financing (AFN) is required for next year?

   USE: (Current Total Assets/Current Sales) x Change in Sales – (Current Liabilities/Current Sales) x  

               Change in Sales – [(NPM x Projected Sales) x (1- Dividend ratio)]. Note that, no dividend is paid.

[Work space]

                        

                  

          Circle A, B, C, or D.                                                                                                          

A. -$ 912

B. -$   723

C.   $ 967

D.   $1,698

3. Assuming you have calculated a negative (less than zero) Additional Funds Needed (AFN) from the

      above data (Table 1) then, and as we discussed in class, which of the following is correct:

  1. The company is taking on too less debt
  2. The company is taking on too much debt
  3. The company should revise its discretionary sources of financing
  4. The company should revise its spontaneous sources of financing
  5. The company has a shortfall in cash

.
Answer #1

A. Total assets = 56600*103% = 58298

Option B, $58298

B. Change in sales = 47000 *3% = 1410

(Current Total Assets/Current Sales) x Change in Sales – (Current Liabilities/Current Sales) x Change in Sales – [(NPM x Projected Sales) x (1- Dividend ratio)]

= (56600/47000) x 1410– (6200/47000) x 1410– [(5% x 48410) x (1- 0)]

= -909

Closest is Option A

c. Option c, The company should revise its discretionary sources of financing